The circumference of circle II is 4 feet longer than the circumference of circle I. How many feet longer is the radius of circle II than the radius of circle I?

Answer by stanbon(75887)

The circumference of circle II is 4 feet longer than the circumference of circle I. How many feet longer is the radius of circle II than the radius of circle I?
---
Formula: Circumference = (pi)(diameter) = (pi)(2r)
So radius = Circumference/(2pi)
----
Circumference of circle I = pi*(2r)
radius of Circle I = r = C/(2pi)
------------------------------------------
Circumference of circle II = 2(pi)r + 4
radius of Circle II = R = [2(pi)r+4]/(2pi) = [(pi)r +2]/pi = r + (2/pi)
------
Circle II's radius is [2/pi] ft longer
===========================================
Cheers,
Stan H.
===================